OpenStudy (anonymous):

What's a Convergent series basically..?? Our professor gave us the idea that if the nth term and its neighberhood terms tends to be a finite(non fluctuating) quantity. But on the internet I am getting answers as divergent even if the nth term tends to zero.. since, then the method for checking it is calculating the sum of the series. eg. 1 + 1/2 + 1/3 + 1/4 + 1/5 + ...+ 1/n in this wikipedia claims its divergent.. but according to our professor it should be convergent anyone please help me out.. i am not sure what is the concept of it!!

OpenStudy (anonymous):

The series you mentioned here (harmonic series) is divergent, because the sum of it tends to be infinite. When a series is convergent it converges to a finite quantity.

OpenStudy (anonymous):

A convergent series is a series whose infinite sum adds up to a finite number.

OpenStudy (anonymous):

thanx a lot .. but i mean what should be the factor for determining whether its convergent or not.. is it only summation tending to a finite quantity or nth term and its neighborhood terms tending to a finite quantity i.e limit exists at n->infinite and \[a _{n} is finite\]

OpenStudy (anonymous):

If \[ \lim_{n\to \infty}a_n\neq 0 \]Then we know it is divergent, but we don't know if it is convergent.

OpenStudy (anonymous):

if \[\lim_{n \rightarrow \infty} a _{n}= 5\] or any finite quantity is the series divergent or convergent?

OpenStudy (anonymous):

If it is greater than \(0\) then it is divergent.

OpenStudy (anonymous):

i had the misconception that if limit exists its convergent.. thanx a lot for clearing that up

OpenStudy (anonymous):

Remember that \(c\times \infty=\infty\) when \(c>0\).

OpenStudy (anonymous):

However when \(c=0\) you get an indeterminate form. It could result in something finite or something infinite.

OpenStudy (anonymous):

yep and so if a series is not convergent it must be divergent?

OpenStudy (anonymous):

The \(\times \infty\) comes from the fact that you're adding infinite terms

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

so \[\lim_{n \rightarrow \infty} a _{n } =0\] it should be convergent but the harmonic series i asked in the ques is divergent...

OpenStudy (anonymous):

When did I say that means it is convergent?

OpenStudy (anonymous):

Okay what is \[ 0\times \infty \to? \]

OpenStudy (anonymous):

Does it go to \(\infty\) or to some finite number?

OpenStudy (anonymous):

it could be anything

OpenStudy (anonymous):

0 to infinite right ..??

OpenStudy (anonymous):

If you understand that indeterminate forms means "inconclusive" then why are you struggling with this limit being "inconclusive" when it goes to 0?

OpenStudy (anonymous):

ohk so more generally i should go for checking summation of a series and see if its finite .. to conclude its convergence..?

OpenStudy (anonymous):

i mean summation of a sequence not series

OpenStudy (anonymous):

There are many tests.

OpenStudy (anonymous):

You can compare it to another series which you know is convergent or not.

OpenStudy (anonymous):

@wio ohk thanx a lot .. i really appreciate you helping me out ..i'll first consult my notes if problem still persists i'll message you.. bye